A three digit number is chosen at random. What is the probability that either all the digits are odd or all the digits are even? Assume the first digit is not zero and that zero is even.
Answer by edjones(8007)

If the 1st digit is odd then 1/2 * 1/2 = 1/4 all odd.
If the 1st digit is even then 1/2 * 1/2 = 1/4 all even.
Since the 1st digit must be either odd or even but not both then 1/4 is the probability that either all the digits are odd or all the digits are even.